### v1

# Scalaron dark matter dynamics: effects of Higgs non-minimal coupling to gravity

Abstract
One of the key features of the $R^{2}$-gravity is the embedding of a scalar field, scalaron, into the gravity sector. The scalaron interacts with the Standard Model (SM) matter fields through Planck-suppressed couplings. If the scalaron serves as a viable dark matter (DM) candidate, it can account for the lack of evidence of DM interactions beyond gravity in experimental and observational probes to date. The realization of the scalaron, as a cold DM candidate, depends on an induced trilinear interaction with the SM Higgs, despite its suppressed coupling strength. Here, we introduce a Higgs non-minimal coupling to gravity that additionally contributes to the induced trilinear interaction with its existing competing part, originated from the $R^{2}$-gravity. We study the interplay between these two contributions in the early universe, which determines both the initial conditions and evolution of the scalaron, leading to cold DM behavior at a later epoch. The trilinear interaction vanish identically for certain combinations of the Higgs non-minimal coupling and the scalaron mass, thereby setting the scalaron density through misalignment mechanism, similar to axions. Consequently, we find that the observed DM relic density is obtained for a scalaron mass between $1.38$ keV and $0.7$ MeV. The lower limit on the mass stems from the LHC constraints on the Higgs non-minimal coupling, whereas the upper bound arises from INTEGRAL/SPI limits on the excess gamma-ray flux from possible scalaron decays to two photons.
1 Introduction
Being in the realm of Einstein’s theory of gravitation, a non-baryonic, pressure-less and non-luminous component of matter, termed as dark matter (DM), is needed for large-scale structure (LSS) formation of our universe. Till date, gravitational imprints of DM have been substantiated with a plethora of cosmological and astrophysical phenomena. However, interactions other than gravity are yet to be confirmed from experiments and observations in operation. In particular, DM interactions with other Standard Model (SM) particles are usually probed in direct, indirect and collider searches. However, its non-observation in these experimental facilities indicates feeble interaction rate, while challenging the viability of canonical WIMP DM scenario in its simplest possible form Profumo:2017hqp ; Bertone:2004pz .
As a nightmare scenario, DM may not interact with SM particles at all, which does not hinder the gravitational clustering, essential for the LSS formation. It only depends on the DM density during the matter-radiation equality epoch, in the cold DM paradigm. Then, the question comes why there is little or absolutely no interaction between DM and SM particles. In this context, extensions of Einstein’s general relativity (GR), particularly $R^{2}$-gravity ($R$ is the Ricci scalar) can provide an explanation to naturally suppressed interactions of DM with the rest of matter fields Cembranos:2008gj ; Corda:2011aa ; Katsuragawa:2016yir ; Katsuragawa:2017wge ; Shtanov:2021uif . In this case, an extra scalar degree of freedom is hidden inside the gravity sector, in addition to two polarizations of gravitational waves. The scalar field becomes disentangled from the gravity sector, while restoring the usual Einstein-Hilbert action DeFelice:2010aj , through a conformal transformation of the original theory with $R^{2}$ term. Since the scalar field, often termed as scalaron starobinsky , comes from the gravity sector, it couples democratically to all other preexisting matter fields but with Planck-scale suppressed couplings after the transformation. Therefore, the scalaron naturally explains the non-observation of DM-SM interactions in existing experiments till date, subject to its viability as DM candidate.
How can the scalaron density satisfy the observed relic density of DM ? The scalar field dynamics, leading to the present-day DM density is a well studied scenario in the context of axion cosmology Preskill:1982cy ; Abbott:1982af ; Dine:1982ah ; Marsh:2015xka . The axion field is misaligned initially from its potential minima. When the Hubble friction becomes negligible compared to the axion mass, coherent oscillation (i.e. associated quanta are in the zero momentum state) starts about the minima of its potential, approximated to a quadratic form. The energy density of the axion field during coherent oscillation redshifts as pressure-less matter, thereby validating it as a cold DM candidate. The observed DM relic density determines the axion mass, $m\sim\mathcal{O}(\mu\rm eV)$ and the initial field value, $\phi_{i}\sim 10^{10}{~\rm GeV}$, for the misalignment scenario during a radiation-dominated universe.
The misalignment mechanism in the context of scalaron has been discussed in Ref.Cembranos:2008gj , without taking effects of scalaron interactions with the SM particles in determining its dynamics in the early universe. An induced trilinear interaction of the scalaron with the Higgs, fixed by the Higgs quartic self-coupling and the vacuum expectation value (VEV), determines the initial condition and the subsequent scalaron dynamics, despite the Planck suppression stemming from its gravitational origin. Thus, the evolution of the scalaron is connected to the Electroweak phase transition (EWPT) Shtanov:2021uif ; Shtanov:2022xew , during which the Higgs field moves from a trivial minima ($\langle h\rangle=0$) to its SM VEV. In this scenario, the coherent oscillation of the scalaron starts only when the Higgs field acquires its SM VEV, unlike the former case, where the Hubble friction decides the onset of the oscillation. As a result, the scalaron saturates the DM density only for a very restrictive mass around $\mathcal{O}$(meV). In contrast, the scalaron mass for the case in Ref.Cembranos:2008gj , varies from $\mathcal{O}$(meV) to $\mathcal{O}$(MeV), depending on the initial field value, which is set independently of Higgs dynamics during the EWPT.
In the present work, we introduce Higgs non-minimal coupling to gravity in addition to the $R^{2}$ term in the Einstein-Hilbert action, for which the above-mentioned scenarios can be realized in a general framework. In particular, the trilinear scalaron-Higgs interaction now gets competing contributions from both $R^{2}$ term and the Higgs non-minimal coupling to gravity. We find that an interplay between these terms decides the initial condition and the scalaron dynamics in the early universe. In fact, the interaction can identically vanish for different combinations of non-minimal coupling and the scalaron mass. Consequently, the scalaron density is set via misalignment dynamics, as in the axion cosmology. In this case, we find that the scalaron mass varies from $\mathcal{O}$(keV) nearly to $\mathcal{O}$(MeV), while satisfying the DM relic density constraint, Higgs global signal strength measurement at Large Hadron Collider (LHC) and $16$ years of gamma-ray observations by INTEGRAL/SPI telescope. For a non-zero scalaron-Higgs interaction, we find that the scalaron mass is constrained to $\mathcal{O}$(meV), irrespective of Higgs non-minimal coupling to gravity permitted by LHC constraints.
The organization of the paper is as follows : the general framework for scalaron dynamics in presence of the Higgs non-minimal coupling has been discussed in Sec.2, followed by the detailed analytic calculation of scalaron dynamics in Sec.3. Then we discuss observational constraints on the non-minimal coupling from LHC experiments and the gamma ray constraints on scalaron decays to two photons from the INTEGRAL/SPI telescope in Sec.4. Finally we summarize our findings in Sec.5.
2 The general framework
We first write the Einstein-Hilbert action with an additional $R^{2}$ term in the Jordan frame as the following starobinsky ; Shtanov:2021uif :
$S_{g}=-\frac{M_{P}^{2}}{3}\int d^{4}x\sqrt{-g}\left(R-\frac{R^{2}}{6m^{2}}\right)\,$, (1)
where $M_{P}(=\sqrt{3/16\pi G}\approx 3\times 10^{18}$ GeV) is the normalized Planck mass. [^1] We have followed the notation of Shtanov:2021uif for the normalized Planck mass $M_{P}$ so that the conformal factor comes out to be $e^{-\frac{\phi}{M_{p}}}$, instead of $e^{-\sqrt{\frac{2}{3}}\frac{\phi}{M_{P}}}$. We do not include the cosmological constant in the gravity lagrangian, since our present study is related to a standard radiation-dominated universe only. At this point, we note that the scale $m$ is going to determine the mass of the scalar field responsible for the DM, and it is in the range of $\rm meV$ to $\rm MeV$.
Now, we assume that the preexisting matter component follows the SM of particle physics and all SM particles other than Higgs, minimally couple to gravity. Higgs non-minimal coupling to gravity has been extensively discussed in the context of Higgs-inflation Bezrukov:2007ep ; Giudice:2010ka ; Rubio:2018ogq . We only write the Higgs part of the matter sector with its non-minimal coupling to gravity as,
$S_{M}\supset\int d^{4}x\sqrt{-g}\left[g^{\mu\nu}\left(D_{\mu}\Phi\right)^{\dagger}D_{\nu}\Phi-\lambda\left(\Phi^{\dagger}\Phi-\frac{v^{2}}{2}\right)^{2}-\xi R\left(\Phi^{\dagger}\Phi-\frac{v^{2}}{2}\right)\right]\,$, (2)
where $\Phi$ is the Higgs doublet, $v~(=246~{\rm GeV})$ is the Higgs VEV at zero temperature and $\lambda$ is the Higgs quartic self-coupling with its SM value, $\lambda\approx 0.13$. We shall demonstrate that the current theory with Eq. (1) and Eq. (2) will generate an interaction between the scalaron and the Higgs at the renormalizable level. In the context of inflation, such theories with both $R^{2}$-gravity and Higgs non-minimal coupling terms have been considered in the literature Ema:2017rqn ; Gorbunov:2018llf ; He:2018mgb ; He:2018gyf ; Gundhi:2018wyz ; Cheong:2021vdb .
To note, the non-minimal coupling, $\xi$ is a dimensionless quantity, which couples Higgs to gravity at the renormalizable level. This coupling can be constrained (detailed in Sec.4) from the measurement of Higgs global signal strength at LHC, given by $|\xi|<1.38\times 10^{15}$ at $95\%$ confidence level (C.L.). The sign of $\xi$ can not be determined uniquely from the signal strength, since only the magnitude of $\xi$ appears in the amplitudes of physical processes, measured at the LHC. However, to ensure the vacuum stability of Higgs potential during high scale inflationary epoch, $\xi$ turns out to be a positive quantity, particularly $\xi\gtrsim 0.06$ Herranen:2014cua . Also for post-inflationary epoch $\xi$ is reported to be positive Figueroa:2017slm . Although we are not concerned with inflationary dynamics in the present context, we work with positive values of $\xi$, i.e. $0\leq\xi\leq 1.38\times 10^{15}$.
To express the total action, (i.e. $S=S_{g}+S_{M}$) in the Einstein frame, we first introduce an auxiliary field, $\sigma$ and write an equivalent action (well known technique in $f(R)$ gravity theories, see for example Refs.Faraoni:2004pi ; DeFelice:2010aj ) as follows :
$S=-\frac{M^{2}_{P}}{3}\int d^{4}x\sqrt{-g}\left[\left(1-\frac{\sigma}{3m^{2}}+\frac{3\xi}{M^{2}_{P}}\left(\Phi^{\dagger}\Phi-\frac{v^{2}}{2}\right)\right)R+\frac{\sigma}{6m^{2}}\right]$ (3)
$+\int d^{4}x\sqrt{-g}\left[g^{\mu\nu}\left(D_{\mu}\Phi\right)^{\dagger}D_{\nu}\Phi-\lambda\left(\Phi^{\dagger}\Phi-\frac{v^{2}}{2}\right)^{2}\right].$
The putative scalar degree of freedom ($\phi$) becomes disentangled from the gravity sector, restoring usual Einstein-Hilbert term via the following conformal transformation Ema:2023dxm :
$g_{\mu\nu}(x^{\mu})=\Omega^{-2}\tilde{g}_{\mu\nu}(x^{\mu}),\qquad\text{with}\quad\Omega=e^{\phi/2M_{P}}=\left(1-\frac{\sigma}{3m^{2}}+\frac{3\xi}{M^{2}_{P}}\left(\Phi^{\dagger}\Phi-\frac{v^{2}}{2}\right)\right)^{1/2}.$ (4)
With this transformation the total action in the Einstein frame becomes,
[…]
$+\int d^{4}x\sqrt{-\tilde{g}}\,\tilde{g}^{\mu\nu}\left[\left(D_{\mu}\tilde{\Phi}\right)^{\dagger}D_{\nu}\tilde{\Phi}+\frac{1}{2M_{P}}\partial_{\mu}\left(\tilde{\Phi}^{\dagger}\tilde{\Phi}\right)\partial_{\nu}\phi+\frac{1}{4M_{P}^{2}}\tilde{\Phi}^{\dagger}\tilde{\Phi}\,\partial_{\mu}\phi\partial_{\nu}\phi\right].$ (5)
[…]
$V(\phi,\tilde{\Phi})\simeq\frac{1}{2}m^{2}M_{P}^{2}\left[1-e^{-\phi/M_{P}}-\frac{3\xi}{M_{P}^{2}}\left(\tilde{\Phi}^{\dagger}\tilde{\Phi}-e^{-\phi/M_{P}}\frac{v^{2}}{2}\right)\right]^{2}+\lambda\left(\tilde{\Phi}^{\dagger}\tilde{\Phi}-e^{-\phi/M_{P}}\frac{v^{2}}{2}\right)^{2},$ (6)
while dropping non-renormalizable terms suppressed by the Planck mass. For sub-Planckian value of $\phi$, i.e. $\phi/M_{P}\ll 1$ and taking the Higgs doublet in the unitary gauge as $\tilde{\Phi}^{T}=(0~~h/\sqrt{2})$, the potential with dominant contributions becomes,
$V(\phi,h)\simeq\frac{1}{2}m^{2}\phi^{2}+\frac{\lambda}{4}\left(h^{2}-v^{2}\right)^{2}+\frac{(h^{2}-v^{2})~\phi}{2M_{P}}\left(\lambda v^{2}-3\xi m^{2}\right).$ (7)
To note, the trilinear coupling between the Higgs and the scalaron, induced after the transformation, receives two distinct contributions - one from $R^{2}$ term and other from the Higgs non-minimal coupling. The interplay between these two terms, as we shall see, will determine the scalaron dynamics in the early universe.
We also incorporate finite temperature correction to the Higgs potential in our scalaron-Higgs effective theory, given as
$\mathcal{L}_{\text{eff}}\supset\frac{1}{2}(\partial_{\mu}\phi)(\partial^{\mu}\phi)+\frac{1}{2}(\partial_{\mu}h)(\partial^{\mu}h)-V(\phi,h)-\frac{1}{6}T^{2}h^{2},$ (8)
where the last term depending on the SM bath temperature ($T$) denotes a finite temperature contribution to the Higgs potential, as approximated in Ref. Rubakov:2017xzr . Within this framework, the scalaron dynamics is studied in conjunction with the Higgs during the EWPT, which sets the initial condition for the scalaron Shtanov:2021uif . Therefore, we first find the extrema of the potential via calculating $dV/d\phi=0$ and $dV/dh=0$, which yields,
$\phi_{0}=\frac{h^{2}-v^{2}}{2m^{2}~M_{P}}\left(3\xi m^{2}-\lambda v^{2}\right),$ (9)
$h_{0}=0~\text{or}~h_{0}^{2}=v^{2}+\frac{\phi}{\lambda M_{P}}\left(3\xi m^{2}-\lambda v^{2}\right)-\frac{T^{2}}{3\lambda}.$ (10)
The non-trivial Higgs minima is modified owing to scalaron-Higgs interaction. For sub-Planckian values of $\phi$, the Higgs minima can well be approximated to its SM value, evident from the following :
$h_{0}^{2}=v^{2}\left(1-\frac{T^{2}}{T^{2}_{c}}\right),~~T_{c}=\sqrt{3\lambda}v\left[1-\frac{\left(3\xi m^{2}-\lambda v^{2}\right)^{2}}{2m^{2}M^{2}_{P}\lambda}\right]^{1/2}\approx\sqrt{3\lambda}v,$ (11)
where $T_{c}$ is the critical temperature of the EWPT within the SM. For $T>T_{c}$, the minima of the Higgs field is at $h_{0}=0$ and for $T<T_{c}$, the minima evolves over temperature and finally settles down to its zero temperature VEV. Consequently, the scalaron minima is dictated by the Higgs dynamics, as manifest in the following,
$\phi_{0}(T)=\begin{cases}\frac{v^{2}}{2m^{2}M_{P}}\left(\lambda v^{2}-3\xi m^{2}\right)~~\text{for $T>T_{c}$},\vskip 5.69046pt\\ \frac{v^{2}}{2m^{2}M_{P}}\left(\frac{T}{T_{c}}\right)^{2}\left(\lambda v^{2}-3\xi m^{2}\right)~~\text{for $T\leq T_{c}$}\vskip 5.69046pt,\\ ~0~~\text{for}~3\xi m^{2}=\lambda v^{2}~\text{and for all $T$}.\end{cases}$ (12)
It is apparent from the above conditions that for non-zero trilinear interaction with the Higgs, the minima of the scalaron becomes non-trivial before the EWPT ; subsequently the minima shifts towards zero as $T\ll T_{c}$. Hence, the initial minima of the scalaron determines the energy density of the coherent oscillation of the field around its final minima, i.e. $\phi_{0}=0$. This is a quite departure from the standard misalignment scenario, where the initial condition is usually set away from the minima, which provides the energy of the coherent oscillation. In the present situation, the source of that energy comes from the interaction with the Higgs, especially due to Higgs quartic self-coupling. Thus the scalaron, being in its natural minima at the early epoch, can account for the DM density at a later epoch.
The interaction with Higgs vanishes when $3\xi m^{2}=\lambda v^{2}$, in presence of a non-zero non-minimal coupling of Higgs to gravity. Consequently, the scalaron minima restores to $\phi=0$ for all temperatures, signifying complete decoupling from the Higgs sector from the very beginning. In this case, the DM density is satisfied by the scalaron, if it is misaligned from its minima in the early universe, reminiscent of the axion cosmology.
Before studying these two scenarios, we write the equation of motion (EOM) for both the scalaron and Higgs field as follows: [^2] We are interested in dynamics of zero momentum modes of the scalaron and the Higgs, thereby writing only the EOMs for the homogeneous part.
$\ddot{\phi}+3H\dot{\phi}+m^{2}\left[\phi+\frac{h^{2}-v^{2}}{2m^{2}M_{P}}\left(\lambda v^{2}-3\xi m^{2}\right)\right]=0,$ (13)
$\ddot{h}+3H\dot{h}+\lambda h\left[h^{2}-v^{2}\left(1-\frac{T^{2}}{T^{2}_{c}}\right)+\frac{\phi}{\lambda M_{P}}\left(\lambda v^{2}-3\xi m^{2}\right)\right]=0,$ (14)
where the Hubble parameter, $H=\dot{a}/a$, in which $a$ denotes the scale factor. The dynamics of the Higgs field becomes independent from $\phi$ for $\phi\ll M_{P}$, evident from Eq.14. For $T<T_{c}$, the Higgs field makes a transition from $h=0$ to $h=v$, thus its mass becomes, $m_{H}\sim\sqrt{\lambda}v$. Therefore, the characteristic time scale of Higgs oscillation is approximately, $t_{h}\sim(\lambda v^{2})^{-1/2}$, whereas the scalaron counterpart is $t_{s}\sim m^{-1}$. For $t_{h}\gg t_{s}$, we can approximate the instantaneous value of the Higgs field with its temperature-dependent VEV ($h_{0}$), as shown in Eq.11. This is achieved with $m\ll m_{H}$ Shtanov:2021uif . Also note that, the scalaron is expected to be a DM candidate, its decay to a pair of Higgs particles is conveniently prevented with this condition. Therefore, the EOM of $\phi$ can be solved independently after replacing the Higgs field with its temperature-dependent VEV into Eq.13 which now takes the following form :
$\ddot{\phi}+3H\dot{\phi}+m^{2}\left[\phi-\frac{T^{2}}{T^{2}_{c}}~\frac{\left(\lambda v^{2}-3\xi m^{2}\right)~v^{2}}{2m^{2}M_{P}}\right]=0.$ (15)
As we solve the equation in the radiation-dominated universe from the epoch of the EWPT ($T=T_{c}$), the time derivatives are conveniently transformed into temperature derivatives, using the relation, $\dot{T}=-HT$, where $H=(\frac{\pi^{2}}{60}g_{*})^{1/2}\frac{T^{2}}{M_{P}}$. Here, we neglect the variation of relativistic degrees of freedom ($g_{*}$). We write the EOM in terms of rescaled dimensionless variables, $\tilde{\phi}=\frac{\phi}{\phi_{0}(T_{c})}$ (for $\phi_{0}\neq 0$) and $\tilde{T}=\frac{T}{T_{c}}$, i.e.
$\tilde{\phi}^{\prime\prime}+\omega^{2}(\tilde{T})\left(\tilde{\phi}-\tilde{T}^{2}\right)=0,$ (16)
where $\phi_{0}(T_{c})=\frac{v^{2}}{2m^{2}M_{P}}\left(\lambda v^{2}-3\xi m^{2}\right)$ and ${\omega^{2}(\tilde{T})=\frac{m^{2}}{H(\tilde{T})^{2}\tilde{T}^{2}}}$. In case of $3\xi m^{2}=\lambda v^{2}$, the EOM simplifies to
$\displaystyle{\bar{\phi}^{\prime\prime}+\omega^{2}(\bar{T})\bar{\phi}=0},$ (17)
but with an arbitrary initial condition away from its minima, i.e. $\phi=\phi_{i}$ at $T=T_{i}$, for which we define $\bar{\phi}=\phi/\phi_{i}$ and $\bar{T}=T/T_{i}$. Notably, the initial temperature in this scenario can be arbitrary; whereas in the former one, the initial temperature is the critical temperature of the EWPT. This is due to apparent decoupling of the scalaron from the Higgs sector, leaving scalaron dynamics unaffected by the phase transition.
3 Scalaron dynamics : effects of Higgs interactions
With the general setup explained earlier, we are all set to discuss two scenarios : in one of which the scalaron dynamics is governed by the Higgs quartic self interaction ($\lambda$) and in the other one, the non-minimal coupling to gravity ($\xi$) decides the dynamics along with an independent initial condition for the scalaron field. We solve Eq.16 and Eq.17 to demonstrate different dynamics at play, leading the scalaron to be a viable DM candidate. Consequently, the scalaron mass required to saturate the observed dark matter relic density differs widely, depending on the underlying dynamics.
3.1 Scenario I : $3\xi m^{2}\neq\lambda v^{2}$
In this scenario, the scalaron stays at its natural minima (as in Eq.12) before the EWPT and only starts moving as $T\leq T_{c}$. The scalaron, initially coupled to the SM thermal bath via the trilinear coupling, eventually decouples. Then it oscillates coherently as a matter field at temperatures well below $T_{c}$. This is a sheer contrast to axion misalignment scenario, in which the oscillation takes place depending only on the ratio of the Hubble parameter and the axion mass. Here, the onset of the oscillation is governed by the scalaron-Higgs interaction in addition to the Hubble to mass ratio. We demonstrate salient features of scalaron dynamics by solving Eq.16 semi-analytically with another change of variable, i.e. $\chi=\tilde{\phi}-\tilde{T}^{2}$, which yields,
$\chi^{\prime\prime}+\omega^{2}(\tilde{T})\left(\chi+\frac{2}{\omega^{2}(\tilde{T})}\right)\,=0.$ (18)
The equation can be solved analytically in the regime $\tilde{T}\leq 1$, for a constant $g_{*}(\approx 100)$, which sets ${\omega^{2}(\tilde{T})=\frac{3m^{2}M_{P}^{2}}{5\pi^{2}T_{c}^{4}\tilde{T}^{6}}}$. Initial conditions are set as : $\chi(\tilde{T}=1)=0$, $\chi^{\prime}(\tilde{T}=1)=-2$, for $\tilde{\phi}(\tilde{T}=1)=1$ and $\tilde{\phi^{\prime}}(\tilde{T}=1)=0$. The above equation further simplifies to a homogeneous one, as $\omega^{2}>>2$ for $\tilde{T}\leq 1$. Hence, the solution of $\chi$ is given by,
[…]
$A$ $=\frac{4}{\sqrt{\alpha}}\Bigg[J_{-\frac{5}{4}}\left(\frac{\sqrt{\alpha}}{2}\right)J_{\frac{1}{4}}\left(\frac{\sqrt{\alpha}}{2}\right)+J_{-\frac{1}{4}}\left(\frac{\sqrt{\alpha}}{2}\right)J_{\frac{5}{4}}\left(\frac{\sqrt{\alpha}}{2}\right)$
$\quad-J_{\frac{1}{4}}\left(\frac{\sqrt{\alpha}}{2}\right)J_{\frac{3}{4}}\left(\frac{\sqrt{\alpha}}{2}\right)-J_{-\frac{1}{4}}\left(\frac{\sqrt{\alpha}}{2}\right)J_{-\frac{3}{4}}\left(\frac{\sqrt{\alpha}}{2}\right)\Bigg]^{-1},$ (20)
where $J_{n}(x)$ denotes the Bessel function of the first kind of order $n$, and ${\alpha=\frac{3m^{2}M_{P}^{2}}{5\pi^{2}T_{c}^{4}}}$.
Figure 1: Scenario-I: Evolution of the scalaron energy density (shown by the pink solid line) in the absence of non-minimal Higgs coupling to gravity. It first redshifts as radiation density (shown by the black dashed line), then as matter density (shown by the blue solid line). This feature is retained even with non-zero non-minimal Higgs coupling, as long as $3\xi m^{2}\neq\lambda v^{2}$.
Although the solution of the original field, $\tilde{\phi}$ can be written using the above expression, the solution of $\chi$ gives important insights about the scalaron as a valid DM candidate. For $\tilde{T}<1$, the solution of $\chi$ can be approximated as, $\chi\sim A\tilde{T}^{3/2}\cos\left({\frac{\sqrt{\alpha}}{2\tilde{T}^{2}}}-\frac{3\pi}{8}\right)$, which in turn sets its (scaled) energy density as the following:
$\rho_{\chi}=\frac{1}{2}\dot{\chi}^{2}+\frac{1}{2}m^{2}\chi^{2}\sim\text{const}\cdot~\tilde{T}^{3}\propto\left(\frac{a}{a_{c}}\right)^{-3}.$ (21)
Here $a_{c}$ is the scale factor corresponding to $T_{c}$. We note that the energy density of $\chi$ redshifts as matter throughout the cosmic history, as shown in Fig.1 (blue solid line). The scalaron can be a viable DM candidate if its density follows similar trend as of $\rho_{\chi}$. We show the energy density of the original field, $\phi$ (depicted by the pink line) in Fig.1. Evidently, it does not follow $\rho_{\chi}$ during initial epochs, instead it follows a radiation-like behavior, apparent from the scaling nature shown by the black-dashed line. Such a behavior can be explained from the solution of $\tilde{\phi}$, given by
$\tilde{\phi}(\tilde{T})=\tilde{T}^{2}+A\sqrt{\tilde{T}}\left[J_{\frac{1}{4}}\left(\frac{\sqrt{\alpha}}{2}\right)J_{-\frac{1}{4}}\left(\frac{\sqrt{\alpha}}{2\tilde{T}^{2}}\right)-J_{-\frac{1}{4}}\left(\frac{\sqrt{\alpha}}{2}\right)J_{\frac{1}{4}}\left(\frac{\sqrt{\alpha}}{2\tilde{T}^{2}}\right)\right].$ (22)
Noticeably, around $T\lesssim T_{c}~(=154~{\rm GeV})$, $\tilde{\phi}(\tilde{T})\sim\tilde{T}^{2}$, in turn $\rho_{\phi}\sim T^{4}$, while the second term in the equation being subdominant. This is a consequence of the scalaron-Higgs interaction, however small, keeps the scalaron coupled to the SM radiation bath for a significant duration of its evolution, depending on the scalaron mass in case of $\xi=0$. For $T<<T_{c}$, eventually the second term dominates, consequently the radiation-to-matter transition of the scalaron density takes place at temperature around $T_{d}=60$ MeV for $m=3.6\times 10^{-3}$ eV (obtained in Eq.24 to saturate present day DM density, discussed later), as shown in Fig.1. This also marks the decoupling of the scalaron from the Higgs, as the Higgs potential takes its usual $T=0$ form. Until the decoupling, the local minima of the scalaron shifts continuously and the scalaron tracks its time-dependent minima, with a characteristic oscillatory feature, apparent from Eq.22 and Eq.12. This is further depicted in Fig.2, by showing the evolution of $\phi$ (pink solid line) and its ambient minima (blue dashed line) at earlier epochs before the scalaron decoupling. We emphasize that the oscillatory feature shown in Fig.2 is inherent to scalaron dynamics since the initial epoch, unlike the standard axion scenarios.
Since, the scalaron energy density at late times, i.e. $T<<T_{c}$ matches with that of $\chi$, we can express the scalaron relic density in terms of the (scaled) energy density of $\chi$ at $T=T_{c}$ as
Figure 2: Scenario-I: Evolutions of the scalaron field (pink solid line) and its minima at early epochs. To note, the scalaron follows its minima that shifts continuously (depicted in blue dashed line) due to induced trilinear interaction with the SM Higgs.
$\Omega_{\phi}=\frac{\rho_{\phi}(T_{0})}{\rho_{c,0}}=\frac{\phi_{0}^{2}(T_{c})\rho_{\chi}(T_{0})}{\rho_{c,0}}=\frac{\phi_{0}^{2}(T_{c})~\rho_{\chi}(T_{c})}{\rho_{c,0}}\left(\frac{T_{0}}{T_{c}}\right)^{3}\frac{g_{*S}(T_{0})}{g_{*S}(T_{c})},$ (23)
where $\rho_{c,0}(\approx 8.2\times 10^{-47}h^{-2}{~\rm GeV^{4}})$ and $T_{0}$ are the critical energy density and the photon temperature of the present universe, while $g_{*S}(T_{i})$ corresponds to the effective number of degrees of freedom contributing to the entropy density of the universe at $T=T_{i}$. In passing, we note that the decoupling temperature of the scalaron does not explicitly enter into the above expression; however, the decoupling must happen before the matter-radiation equality epoch, i.e. $T\sim 1$ eV, necessary condition for successful structure formation. In the present context, the decoupling happens around $60$ MeV, for which the scalaron density trivially satisfies the Big Bang Nucleosynthesis bound Fields:2019pfx on any extra radiation density present in the cosmic soup.
In calculating the relic density, the initial kinetic energy of $\chi$ only contributes, since at $T=T_{c}$, $\chi=0$, which renders $\rho_{\chi}(T_{c})=\dot{\chi}^{2}(T_{c})/2=2H^{2}_{c}\approx 2.05\times 10^{-27}~\rm{GeV^{2}}$. For $\xi=0$, $\phi_{0}(T_{c})=\frac{\lambda v^{4}}{2m^{2}M_{P}}=7.9\times 10^{-11}{\rm GeV}\left(\frac{{~\rm GeV}}{m}\right)^{2}$, and the observed DM relic density Planck:2018vyg is set by the scalaron mass only, i.e.
$\Omega_{\phi}h^{2}=0.12\left(\frac{3.6\times 10^{-12}{~\rm GeV}}{m}\right)^{4},$ (24)
where we have used $g_{*S}=3.91$ and $g_{*S}(T_{c})=g_{*}(T_{c})=100$. This reproduces the primary result of Refs.Shtanov:2021uif ; Shtanov:2024nmf ; Shtanov:2025nue within our general treatment of scalaron dynamics.
For non-zero values of $\xi$, particularly $\xi\ll\frac{\lambda v^{2}}{3m^{2}}$, the scalaron can in principle saturate the observed DM density only via the non-minimal Higgs coupling to gravity, as $\phi_{0}(T_{c})\approx 3\xi v^{2}/2M_{P}$. However, this would require an enormously large value of $\xi$ as shown in the following, which has been already excluded by the LHC constraints, as indicated earlier :
$\Omega_{\phi}h^{2}=0.12\left(\frac{\xi}{1.95\times 10^{26}}\right)^{2}.$ (25)
Therefore, the scalaron dynamics for $3\xi m^{2}\neq\lambda v^{2}$, even with non-zero allowed values of $\xi$, is completely controlled by the Higgs quartic self-coupling. In fact, the initial energy density of the scalaron depends on $\phi_{0}$ and the Hubble parameter at $T=T_{c}$, which are determined by the SM value of $\lambda$. As a result, the scalaron mass is restricted to a value of $3.6\times 10^{-3}$ eV, to satisfy the DM relic density constraint. This restriction on the scalaron mass is lifted when the trilinear interaction becomes either sufficiently small or identically vanishes for $\lambda v^{2}=3\xi m^{2}$, which is studied next.
Figure 3: Scenario-II: Evolution of the scalaron field ($\phi/\phi_{i}$) (left panel) and its energy density ($\rho_{\phi}$) (right panel) for $\xi\sim\mathcal{O}(10^{15})$, which fixes $m\sim\mathcal{O}$(keV). The initial field value $\phi_{i}$, is determined from the observed DM relic density. See text for details.
3.2 Scenario II : $3\xi m^{2}=\lambda v^{2}$
In this scenario, the trilinear interaction between Higgs and the scalaron can identically vanish for a wide range of allowed values of $\xi$, thereby the scalaron remains decoupled from the SM thermal bath at the outset of the evolution. Unlike the previous case, the scalaron density remains trivial if the scalaron stays at its minima, $\phi_{0}=0$. Therefore, the scalaron has to be misaligned from its minima with nearly zero velocity ($\dot{\phi}=0$), which facilitates coherent oscillation of the field around its trivial minima at a later time. This is apparent from the solution of Eq.17, i.e.
$\frac{\phi(\bar{T})}{\phi_{i}}=-\frac{\pi\sqrt{\alpha\bar{T}}}{2\sqrt{2}}\left[J_{\frac{5}{4}}\left(\frac{\sqrt{\alpha}}{2}\right)J_{-\frac{1}{4}}\left(\frac{\sqrt{\alpha}}{2\bar{T}^{2}}\right)+J_{-\frac{5}{4}}\left(\frac{\sqrt{\alpha}}{2}\right)J_{\frac{1}{4}}\left(\frac{\sqrt{\alpha}}{2\bar{T}^{2}}\right)\right],$ (26)
where ${\alpha=\frac{3m^{2}M_{P}^{2}}{5\pi^{2}T_{i}^{4}}}$ and $\phi_{i}$ being the (misaligned) field value at the initial epcoh, $T=T_{i}$. For $\bar{T}<1$, the solution is approximated again as, ${\phi\sim\phi_{i}\bar{T}^{3/2}\cos\left({\frac{\sqrt{\alpha}}{2\bar{T}^{2}}-\frac{3\pi}{8}}\right)}$, which ensures the cold DM behavior of the scalaron density from the onset of the coherent oscillation, unlike the previous scenario, in which there is at first radiation-like behavior, followed by the matter-like one. Here, as long as the Hubble friction is dominant, $\rho_{\phi}$ remains constant at its initial value, $\rho_{\phi}(T_{i})=\frac{1}{2}m^{2}\phi^{2}_{i}$. The density starts decreasing once the Hubble friction becomes comparable with the scalaron mass and subsequently the scalaron oscillates coherently. The temperature at which coherent oscillation starts is estimated from $3H\sim m$ is given by,
$T_{osc}\approx\left(\frac{mM_{P}}{3}\right)^{1/2}\left(\frac{60}{\pi^{2}g_{*}(T_{osc})}\right)^{1/4}.$ (27)
In Fig.3, evolutions of the scalaron field (left panel) and its density (right panel) have been shown for a typical scalaron mass, $m=1$ keV. It is apparent that the oscillation starts around, $T_{osc}=4.96\times 10^{5}$ GeV, in agreement with Eq.27. As $\rho_{\phi}$ is approximately constant from $T_{i}=10^{7}$ GeV until $T=T_{osc}$, i.e. $\rho_{\phi}(T_{osc})\approx\rho_{\phi}(T_{i})$. So, the relic density can be expressed in terms of $\phi_{i}$ and $m$, utilizing Eq.27, as follows :
$\Omega_{\phi}=\frac{\rho_{\phi}(T_{i})}{\rho_{c,0}}\left(\frac{T_{0}}{T_{osc}}\right)^{3}\frac{g_{*S}(T_{0})}{g_{*S}(T_{osc})}\implies\Omega_{\phi}h^{2}\approx 0.12\left(\frac{m}{10^{-6}{~\rm GeV}}\right)^{1/2}\left(\frac{\phi_{i}}{7\times 10^{10}{~\rm GeV}}\right)^{2},$ (28)
where we have used $g_{*S}(T_{osc})=g_{*}(T_{osc})=100$. We can now choose $\xi$, which fixes the scalaron mass through the relation, $3\xi m^{2}=\lambda v^{2}$, since all other parameters are fixed at their SM value. Consequently, the remaining free parameter, $\phi_{i}$, is determined from the observed DM relic density. Unlike the previous scenario, the scalaron mass can vary depending on $\xi$, that can vary as high as $\mathcal{O}(10^{15})$. In the left panel of Fig.4, we depict the allowed parameter space of $m_{\phi}$ (and corresponding $\xi$ value) and $\phi_{i}$, for which the DM relic density is satisfied. In particular, the constant contour (blue solid line) of $\Omega_{DM}h^{2}=0.12$ has been shown depicting the central value of the DM relic density parameter. The region above the blue solid line in Fig.4 (left panel) is excluded, as the scalaron would be over-abundant, while the lower region corresponds to the scalaron contributing only a fraction of the DM abundance.
We note that the trilinear scalaron-higgs interaction vanishes only when the Higgs non-minimal coupling is introduced, thereby expanding the parameter space of the scalaron mass. Unlike scenario-I, here we have an extra freedom of choosing $\phi_{i}$ that opens a new parameter space for the scalaron mass to be probed in the near future.
Figure 4: Scalaron as dark matter in scenario-II : (Left panel) Constraints on the scalaron mass and initial field value from observed DM relic density, Higgs global signal strength and INTEGRAL/SPI limits on the excess photon flux from the scalaron decay to two photons. (Right panel) The scalaron decay rate to two photons and its exclusion limit at $95\%$ C.L., taken from Ref.Calore:2022pks . See text for details.
4 Constraints
In this section, we discuss both upper and lower limits of the scalaron mass for scenario-II, which are not solely determined from the observed DM relic density. The allowed scalaron mass, as we shall elucidate in the following subsections, varies from $1.38$ keV to $0.7$ MeV (see Fig.4). The lower limit of the scalaron mass is directly related to the upper bound on Higgs non-minimal coupling to gravity from LHC measurements via the relation $m_{\rm min}=(\lambda v^{2}/3\xi_{\rm max})^{1/2}$. Besides, the scalaron mediates an attractive Yukawa-like force between two non-relativistic particles, which is constrained from the torsion-balance experiment. The non-observation of the fifth force puts the lower limit on the scalaron mass as $2.7$ meV Kapner:2006si ; Adelberger:2006dh ; Cembranos:2008gj , which becomes redundant in the presence of LHC constraints on $\xi$ that sets $m_{\rm min}=1.38$ keV. The upper bound originates from the non-observation of an excess photon flux at INTEGRAL/SPI telescope, assuming the scalaron DM decaying to a pair of photons at the current epoch within our galaxy.
4.1 LHC Constraints on $\xi$
As already explained, the role of Higgs non-minimal coupling is pivotal in deciding the scalaron dynamics; however the bound on $\xi$ in the non-inflationary scenario is not well studied. Here we show the upper bound on $|\xi|$ from the measurement of Higgs global signal strength by LHC, following Ref.Atkins:2012yn . This bound is derived when the contribution from $R^{2}$-gravity becomes sub-dominant at $T\approx 0$ universe. Hence, the kinetic term of the SM Higgs is affected only by the non-minimal coupling, $\xi$, which requires the redefinition of the original Higgs field. As its result, Higgs couplings to all other SM particles are suppressed by a factor $\kappa$ Bezrukov:2007ep ; Atkins:2012yn given by,
$\kappa\approx\left(1+\frac{9\xi^{2}v^{2}}{M_{P}^{2}}\right)^{-1/2},$ (29)
while restoring canonical form of its kinetic term. The modification in Higgs couplings to other SM fields, in principle has imprints on the Higgs signal strength measurement at LHC, which is a standard practice to quantify any possible deviation from the SM prediction. In a narrow-width approximation of Higgs production cross-section and its decay, the Higgs signal strength of a process, $ii\rightarrow h\rightarrow ff$ is given by,
$\mu_{if}=\frac{\sigma(ii\rightarrow h)\cdot\mathrm{BR}(h\rightarrow ff)}{\sigma_{\mathrm{SM}}(ii\rightarrow h)\cdot\mathrm{BR}_{\mathrm{SM}}(h\rightarrow ff)}=\kappa^{2}\,$, (30)
where the deviation from the SM depends on $\kappa^{2}$ in our scenario. We use the global Higgs signal strength data, measured by both ATLAS ($\mu=1.05\pm 0.06$ ATLAS:2022vkf ) and CMS ( $\mu=1.002\pm 0.057$ CMS:2022dwd ) to estimate the upper bound on $\kappa^{2}$, and in turn on $|\xi|$. Combining these results, we find the upper limit on $|\xi|$ to be $1.25\times 10^{15}$ at the $95\%$ C.L., which results into the lower bound of the scalaron mass to be $1.38$ keV in scenario-II. This also explains why the scalaron density can not saturate the DM relic density in the scenario-I only with the Higgs non-minimal coupling, as shown in Eq.25.
4.2 Constraints on scalaron decay to photons
The viability of the scalaron as a DM candidate hinges on its lifetime to be comparable to the age of the universe. Since the scalaron universally couples to the SM matter fields, it can decay to most of the SM particles, if allowed by kinematics. To note, for $m<2m_{e}$, $m_{e}$ being the electron mass, the only possible decay is to a pair of photons, whose rate should be such that the scalaron becomes cosmologically stable. Although the scalaron does not couple to massless gauge fields at the tree-level due to Weyl invariance, [^3] The photon action, $\int d^{4}x\sqrt{-g}~F_{\mu\nu}F^{\mu\nu}$ is invariant under trnasformation defined in Eq.4 the coupling is generated at one loop-level, primarily due to interactions with fermions and massive gauge bosons, as follows Goldberger:2007zk ; Cembranos:2008gj ; Katsuragawa:2017wge :
$\mathcal{L}_{\phi-F}=\frac{\phi}{2M_{P}}\frac{\alpha_{EM}c_{EM}}{8\pi}F_{\mu\nu}F^{\mu\nu},$ (31)
where $\alpha_{EM}(=1/137)$ is the electromagnetic fine structure constant and $c_{EM}$ encodes the loop-induced factors. This is reminiscent of the axion–photon coupling, but with a fixed interaction strength stemming from its gravitational origin. Consequently, the decay rate of scalaron to two photons is dependent only on the scalaron mass and given by,
$\Gamma_{\phi\rightarrow\gamma\gamma}=3.8\times 10^{-33}s^{-1}\left(\frac{m}{100~{\rm KeV}}\right)^{3}\,$, (32)
where $c_{EM}=11/3$ Cembranos:2008gj , for scalaron being lighter than the SM fermions and gauge bosons. Evidently, the bound on the lifetime of DM, i.e. $\tau_{DM}\gtrsim 5\times 10^{18}s$ Audren:2014bca is easily satisfied by the scalaron. However, the photon flux produced from scalaron decays can constrain the decay rate from the observation of an excess photon flux over known astrophysical backgrounds. In this case, gamma-ray observations with SPI, a spectrometer telescope aboard the INTEGRAL satellite, operating in the energy range of 18 keV to 8 MeV Winkler:2003nn will be most sensitive. The photon signal from non-relativistic decays of scalaron in the present epoch yields a monochromatic spectrum, peaking at a single energy bin around, $m/2$. Consequently, the scalaron mass is constrained from the INTEGRAL/SPI data, unlike scenarios with axion-like particles (ALPs) with an undetermined U(1) symmetry breaking scale involved Alonso-Alvarez:2019ssa ; Arias:2012az . Nevertheless, there are existing limits on the decay rate of light DM decaying to two photons Calore:2022pks ; Fischer:2022pse , which can be used to constrain our scalaron scenario. In the right panel of Fig.4, we show the exclusion region on the decay rate (green shaded region) using the INTEGRAL/SPI data, taken from Ref.Calore:2022pks . Authors of Ref.Calore:2022pks analyzed 16 years of the INTEGRAL/SPI data, taking the NFW profile of galactic DM density to constrain the decay rate at 95% confidence level. We compare the decay rate of the scalaron (shown in red solid line) as derived in Eq.32 to find the upper bound of the scalaron mass as $m_{\rm max}=0.7$ MeV. This bound is derived with the assumption that the scalaron satisfies total DM density, otherwise the exclusion limits become weaker.
5 Summary and Outlook
To summarize, we have studied the scalaron DM dynamics in presence of scalaron-Higgs trilinear interaction, which consists of Higgs quartic coupling and Higgs non-minimal coupling to gravity. Depending on the relative contributions of these couplings the scalaron dynamics is modified and its mass is significantly altered to account for observed DM density. In particular, we provide a natural explanation for vanishing scalron-Higgs interaction, leading to keV-MeV DM candidate.
We note that the interplay between Higgs quartic coupling and the non-minimal coupling in the light DM context is demonstrated with a possibility of detection in highly sensitive X-ray/gamma ray telescopes. However, there is significant uncertainty over the experimental determination of these couplings, which can affect scalaron dynamics in the early universe. In fact, measuring the Higgs quartic coupling at colliders is challenging, as it requires cross-section measurements of processes involving the three-Higgs vertex, which are highly suppressed Fuks:2017zkg ; Stylianou:2023tgg . On the other hand, the limits on the Higgs non-minimal coupling in low energy effective theory with $R^{2}$ modification to the Einstein’s GR is an unexplored topic, to the best our knowledge. Here, we have used the LHC data to give an upper bound on the non-minimal coupling assuming the $R^{2}$ term subsides for in the late universe. Therefore, the scalaron dynamics and its mass limits can modify with improvements in understanding of these Higgs interaction along with its experimental verifications.
Acknowledgment
SGC acknowledges funding from the Anusandhan National Research Foundation (ANRF), Govt. of India, under the National Post-Doctoral Fellowship (File no.PDF/2023/002066). SGC thanks Ananda Dasgupta for useful discussions. DG acknowledges the Post-doctoral Fellowship from IISER Kolkata for most of the duration of this work, and from SINP Kolkata for the remaining period.